Crew 181 Daily Summary Report
May 19th, 2017
MDRS Daily Summary Report for SOL 6
Summary Title: Foggy Helmets and Ideal Hab
Mission Status: Tiring EVA and Chana Masala
Sol Activity Summary: We were off to a rough start when our early morning EVA yielded less than spectacular results due to foggy helmets (see EVA report). This led me to be a very grumpy commander (and it didn’t help that we stayed up way too late watching A Close Encounter of a Third Kind, resulting on a lack of sleep.) Some of the EVA crew then had their breakfast routine, others went back to bed.
Lunch was prepared by Juan, a simple mac ‘n cheese and peas. After lunch, we had an interesting discussion about what would make up the ideal Hab on Mars. Some things were simple to remark – helmets that don’t fog, working radio comms, better lighting in the science lab, enclosed tunnel system. Others were interesting and quirky – the Hab should have a glue gun, ideally an art exhibition space/gallery, a beehive in the GreenHab… an unlimited supply of Red Lobster cheddar biscuits.
Janet and Charlie worked together to troubleshoot a mechanism on the observatory rotating roof. They are happy to say that it is now functional! Evening EVA was great, no foggy helmets (well, at least not as bad) and Janet, Cassandra, and Olly all got to work on their projects.
Look Ahead Plan: If the weather cooperates, Janet will be using the observatory tonight and everyone else will want a chance to look into the telescope (although Janet will be the only one working it.) We are sad to be saying goodbye to Olly tonight. We are submitting an EVA request for tomorrow.
Anomalies In Work: None at the current time.
Weather: Partly cloudy, lower than average temperature.
Crew Physical Status: All in good spirits.
EVA: Both EVAs today were conducted around the Hab.
Reports to be Filed: journalist report, engineering report, EVA report, and EVA request.
Support Requested: None at the current time.
Prepared by Cassandra
Commander, Crew 181